# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v1

@@ Appendix E Proofs of the Main Results
    $=:\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(\hat{t}_{A})=o_{P_{n}}(1)$
    Note that $\mathbb{E}[W_{i}(t)]=\mathbb{E}[W_{i}(t)|\hat{s}_{A,k}]=0$, and that $\{W_{i}(t)\}_{i\in\mathcal{I}_{j,k}}$ are iid conditional on $\hat{s}_{A,k}$ for any $t$. Hence it’s enough to show that
    $Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(\hat{t}_{A})\right]\leq\sup_{t}Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\right]$
    $=\sup_{t}\mathbb{E}\left[Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\Bigm{|}\hat{s}_{A,k}\right]\right]=\sup_{t}Var\left[W(t)\right]\to 0$
    where the first equality follows from the Law of Total Variance. We have
    $\sup_{t}Var\left[W(t)\right]$
    $\leq\sup_{t}\mathbb{E}\left[\left|\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\right|\right]$
    $\hskip 5.0pt+\sup_{t}\mathbb{E}\left[\left(F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right)^{2}\right]$
    $\hskip 5.0pt+2\sup_{t}\Bigl{|}\mathbb{E}\Bigl{[}\{\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\}$
    ×$\hskip 60.0pt\times\{F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\}\Bigr{]}\Bigr{|}$
    $\leq 4\sup_{t}\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|$
    and $\sup_{t}\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|\to 0$ follows from (uniform) equicontinuity of $P(Y(j)\leq t)$ by A2(i) and the fact that $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii).
  • Step 2: Equation (E.5)
@@
    (i) is a consequence of =$\mathcal{F}=\{f_{t}:\mathbb{R}\rightarrow[0,1],w\mapsto f_{t}(w)=\mathbb{I}(w\leq t):t\in\mathbb{R}\}$ being Donsker and Pre-Gaussian uniformly in $P\in\{P^{*}:(Y-s_{A,P}(X),D,X)\sim P^{*},(Y,D,X)\sim P\in\mathcal{P}\}$; (ii) follows from consistency of M-estimators (e.g., Corollary 3.2.3 in Van Der Vaart and Wellner (2023)), given that $P(Y(j)\leq t)$ is equicontinuous by A2(i) and $t_{A,P}$ is unique and well-separated by A2(iii); (iii) follows from equicontinuity of $P(Y(j)\leq t)$ by A2(i) and since $|\hat{t}_{A}-t_{A,P_{n}}|\overset{P_{n}}{\longrightarrow}0$.
  • Step 3: Equation (E.6)
@@
    $\hskip 15.0pt+\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\sqrt{K}}{\sqrt{n_{1}}}o_{P_{n}}(1)-\frac{\sqrt{K}}{\sqrt{n_{0}}}o_{P_{n}}(1)\right)$ (E.7)
@@
    where (E.7) follows from (E.4) and (E.5).
@@
The inequalities for the one-sided confidence intervals follow directly from consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U})$ and from Theorem E.1, since $\bar{\theta}_{L,P_{n}}\leq\theta_{L,P_{n}}^{*}\leq\theta_{P_{n}}\leq\theta_{U,P_{n}}^{*}\leq\bar{\theta}_{U,P_{n}}$. For the two-sided case, the inequality follows from Theorem E.1 and consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ by applying Proposition 3 in Stoye (2009) centering the estimators at the outer bounds $\bar{\theta}_{L,P_{n}}$ and $\bar{\theta}_{U,P_{n}}$. Since these are in general not sharp, the equality in Stoye (2009) becomes an inequality.
Consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ follows from the weak law of large numbers, equicontinuity of $P(Y(j)\leq t)$ by A2(i), $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii) and $\hat{t}_{A}-t_{A,P_{n}}\overset{P_{n}}{\longrightarrow}0$ (see Step 2 in proof of Theorem E.1). ∎
@@
To show $\sqrt{n}(\bar{\theta}_{A,P_{n}}-\theta_{A,P_{n}}^{*})=o_{P_{n}}(1)$, it’s enough to show $\sqrt{n}\left(\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}\right)=o_{P_{n}}(1)$ since $K$ is fixed. Define $\Delta_{P}(t,s|x)=F_{P,1}(t,s|x)-F_{P,0}(t,s|x)$, where $F_{P,j}(t,s|x)=P(Y(j)\leq s(X)+t|X=x)$ and note that $\Delta_{P}(t,s)=\mathbb{E}[\Delta_{P}(t,s|X)]$. Then, $|\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}|=|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$, since $t_{A,P_{n}}=0$ by definition of $s^{*}_{A,P_{n}}$ (see proof of Theorem 1).
It also holds that $|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|\leq|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$. To see this, focus on the case of the lower bound (the upper bound case is analogous). We have from Theorem 1 that $\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})\geq\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})$, and by definition of $\hat{t}_{L}$ it holds that $\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\geq\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Hence, $|\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})-\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})|=\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\leq\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Therefore, it’s enough to show that $\sqrt{n}|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|=o_{P_{n}}(1)$.
@@ A Taylor expansion gives
$|\Delta_{P_{n}}(0,s|x)-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}}|x)|\lesssim(s(x)-s^{*}_{A,P_{n}}(x))^{2}$ (E.8)
@@
Online Appendix I Relaxing constant propensity score
@@
    Constant propensity score within groups: Let there be fixed $\bar{G}>0$ groups, such that the propensity score is fixed within-group (and bounded away from zero and one). Define $n_{1,g}$ to be the number of observations in group $g$ assigned to treatment and $n_{0,g}$ the number of observations assigned to control. Finally, let $G_{i}\in\{1,\dots,\bar{G}\}$ be the group to which observation $i$ belongs. Then, a new cross-fitting estimator for the lower bound can defined as
@@
    where the data is split such that for each fold $k$ there is balance in the number of treated and control units within each group. The asymptotic distribution follows very similarly to the proof of Theorem 3, by applying the arguments in (E.4) and (E.5) by each group-fold combination, instead of by fold. The asymptotic variance-covariance matrix of the estimators is given by
@@
Online Appendix II Definition of Two-Sided Confidence Intervals
@@
Importantly, calculating $\widehat{CI}_{\alpha}$ relies on choosing the hyperparameter $h_{n}$. A possible choice is Stoye’s suggestion of $h_{n}=n^{-1/2}(\log\log n)^{1/2}$, and other possibilities are $h_{n}=n^{-1/2}(\log n)^{1/2}$ and $h_{n}=n^{-1/2}(Q\log\log n)^{1/2}$ for $Q\geq 2$, as suggested in Andrews and Soares (2010). Discussing a best choice for $h_{n}$ is beyond the scope of this paper, but a practical procedure for applications is to assess the sensitivity of the CI to different choices of $h_{n}$.
Online Appendix III Relaxing Assumption 2(iii)
For the current form of the estimators in Definition 2(iii), some variant of 2(iii) is necessary. For example, if $\Delta_{P}(t,s_{L,P})$ is constant in $t$, then $\widehat{\theta}_{L}$ would converge to the maximum of a Gaussian process.
One way to relax this assumption is by proposing changes in the definition of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$ in order to avoid the optimization problem. For example, $t$ could be fixed at $t=0$, or learned from other folds inside the cross-fitting procedure. When estimating $\hat{s}_{L,k}$, one could estimate $\hat{t}_{k}$ (with data from all except the $k$-th fold) which maximizes the expression inside the $k$-th fold, and estimate the lower bound as $\frac{1}{n}\sum_{i=1}^{n}\left(\frac{D_{i}}{\hat{p}(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})-\frac{1-D_{i}}{1-\hat{p}(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})\right)$. This would be valid since it’s equivalent to a new estimator $\tilde{s}_{L,k}=\hat{s}_{L,k}+\hat{t}_{k}$, and there is no optimization in the final estimator.
Alternatively, the asymptotic distribution of the estimator could be studied without the condition of a unique maximizer, using the fact that $\mathbb{G}_{P}(t)=\sqrt{n}\left(\widehat{\Delta}_{P}(t,\hat{s}_{L,k})-\Delta_{P}(t,s_{L,P})\right)$ converges to a Gaussian process (see proof of Theorem E.1).
Online Appendix IV Additional Tables

### v2

@@ A.5 Proofs of the Main Results
    $=:\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(\hat{t}_{A})$
    $\leq\sup_{t}\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)=o_{P_{n}}(1)$
    The result follows from stochastic equicontinuity of $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)$ plus pointwise convergence in $t$ to zero (see, e.g., Theorem 22.9 in Davidson (2021)).
    Note that
    $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\Bigl{[}\left(\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\right)$
    $\hskip 90.0pt-\left(F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right)\Bigr{]}$
    $=\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-F_{P_{n},j}(t,\hat{s}_{A,k})\right)$
    $\hskip 20.0pt-\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)-F_{P_{n},j}(t,s_{A,P_{n}})\right),$
    and so it is enough to show stochastic equicontinuity of
    $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq\tilde{s}(X_{i})+t)-F_{P_{n},j}(t,\tilde{s})\right)$
    for an arbitrary random variable $\tilde{s}$ independent of $\mathcal{I}_{j,k}$. This follows since =$\mathcal{F}=\{f_{t}:\mathbb{R}\rightarrow[0,1],w\to f_{t}(w)=\mathbb{I}(w\leq t):t\in\mathbb{R}\}$ is Donsker and Pre-Gaussian uniformly in $P\in\{P^{*}:(Y-s(X),D,X)\sim P^{*},(Y,D,X)\sim P^{\prime}\in\mathcal{P},s\in\{s^{\prime}:\mathcal{X}\to\mathcal{Y}\}\}$. This is the case since the covering number of $\mathcal{F}$ for any such $P$ does not depend on $s$.
    Finally, for verifying pointwise convergence, note that
    $\mathbb{E}[W_{i}(t)]=\mathbb{E}[W_{i}(t)|\hat{s}_{A,k}]=0.$
    Hence, it is enough to show that
    $Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\right]$ $=\mathbb{E}\left[Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\Bigm{|}\hat{s}_{A,k}\right]\right]$
    $=\mathbb{E}\left[Var\left[W_{i}(t)\Bigm{|}\hat{s}_{A,k}\right]\right]$
    $\leq Var\left[W(t)\right]\to 0$
    where the first equality follows from the Law of Total Variance, and the second from the fact that $\{W_{i}(t)\}_{i\in\mathcal{I}_{j,k}}$ are iid conditional on $\hat{s}_{A,k}$ for any $t$. We have
    $Var\left[W(t)\right]$
    $\leq\mathbb{E}\left[W(t)^{2}\right]$
    $\leq 2\mathbb{E}\left[\left|\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\right|\right]$
    $\hskip 5.0pt+2\mathbb{E}\left[\left(F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right)^{2}\right]$
    $\leq 4\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|,$
    and $\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|\to 0$ follows from equicontinuity of $P(Y(j)\leq t)$ by A2(i) and the fact that $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii).
  • Step 2: Equation (A.5)
@@
    (i) follows from the same argument for stochastic equicontinuity as in Step 1 (equation A.4). (ii) follows from consistency of M-estimators (e.g., Corollary 3.2.3 in Van Der Vaart and Wellner (2023)), given that $P(Y(j)\leq t)$ is equicontinuous by A2(i) and $t_{A,P}$ is unique and well-separated by A2(iii); (iii) follows from equicontinuity of $P(Y(j)\leq t)$ by A2(i) and since $|\hat{t}_{A}-t_{A,P_{n}}|\overset{P_{n}}{\longrightarrow}0$.
  • Step 3: Equation (A.6)
@@
    $\hskip 15.0pt+\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\sqrt{K}}{\sqrt{n_{1}}}o_{P_{n}}(1)-\frac{\sqrt{K}}{\sqrt{n_{0}}}o_{P_{n}}(1)\right)$ (A.7)
@@
    where (A.7) follows from (A.4) and (A.5).
@@
The inequalities for the one-sided confidence intervals follow directly from consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U})$ and from Theorem A.1, since $\bar{\theta}_{L,P_{n}}\leq\theta_{L,P_{n}}^{*}\leq\theta_{P_{n}}\leq\theta_{U,P_{n}}^{*}\leq\bar{\theta}_{U,P_{n}}$. For the two-sided case, the inequality follows from Theorem A.1 and consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ by applying Proposition 3 in Stoye (2009) centering the estimators at the outer bounds $\bar{\theta}_{L,P_{n}}$ and $\bar{\theta}_{U,P_{n}}$. Since these are in general not sharp, the equality in Stoye (2009) becomes an inequality.
Consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ follows from the weak law of large numbers, equicontinuity of $P(Y(j)\leq t)$ by A2(i), $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii) and $\hat{t}_{A}-t_{A,P_{n}}\overset{P_{n}}{\longrightarrow}0$ (see Step 2 in proof of Theorem A.1). ∎
@@
To show $\sqrt{n}(\bar{\theta}_{A,P_{n}}-\theta_{A,P_{n}}^{*})=o_{P_{n}}(1)$, it is enough to show $\sqrt{n}\left(\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}\right)=o_{P_{n}}(1)$ since $K$ is fixed. Define $\Delta_{P}(t,s|x)=F_{P,1}(t,s|x)-F_{P,0}(t,s|x)$, where $F_{P,j}(t,s|x)=P(Y(j)\leq s(X)+t|X=x)$ and note that $\Delta_{P}(t,s)=\mathbb{E}[\Delta_{P}(t,s|X)]$. Then, $|\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}|=|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$, since $t_{A,P_{n}}=0$ by definition of $s^{*}_{A,P_{n}}$ (see proof of Theorem 1).
It also holds that $|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|\leq|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$. To see this, focus on the case of the lower bound (the upper bound case is analogous). We have from Theorem 1 that $\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})\geq\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})$, and by definition of $\hat{t}_{L}$ it holds that $\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\geq\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Hence, $|\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})-\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})|=\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\leq\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Therefore, it is enough to show that $\sqrt{n}|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|=o_{P_{n}}(1)$.
@@ A Taylor expansion gives
$|\Delta_{P_{n}}(0,s|x)-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}}|x)|\lesssim(s(x)-s^{*}_{A,P_{n}}(x))^{2}$ (A.8)
@@
Online Appendix OA
Supplemental Appendix to “Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach”
OA.1 Relaxing constant propensity score
@@
    Constant propensity score within groups: Let there be fixed $\bar{G}>0$ groups, such that the propensity score is fixed within-group (and bounded away from zero and one). Define $n_{1,g}$ as the number of observations in group $g$ assigned to treatment and $n_{0,g}$ the number of observations assigned to control. Finally, let $G_{i}\in\{1,\dots,\bar{G}\}$ be the group to which observation $i$ belongs. Then, a new cross-fitting estimator for the lower bound can defined as
@@
    where the data is split such that for each fold $k$, there is balance in the number of treated and control units within each group. The asymptotic distribution follows very similarly to the proof of Theorem 3 by applying the arguments in (A.4) and (A.5) by each group-fold combination, instead of by fold. The asymptotic variance-covariance matrix of the estimators is given by
@@
OA.2 Definition of Two-Sided Confidence Intervals
@@
Importantly, calculating $\widehat{CI}_{\alpha}$ relies on choosing the hyperparameter $h_{n}$. A possible choice is Stoye’s suggestion of $h_{n}=n^{-1/2}(\log\log n)^{1/2}$, and other possibilities are $h_{n}=n^{-1/2}(\log n)^{1/2}$ and $h_{n}=n^{-1/2}(Q\log\log n)^{1/2}$ for $Q\geq 2$, as suggested in Andrews and Soares (2010). Discussing the best choice for $h_{n}$ is beyond the scope of this paper, but a practical procedure for applications is to assess the sensitivity of the CI to different choices of $h_{n}$.
OA.3 Relaxing Assumption 2(iii)
I propose two alternatives to Assumption 2(iii). The first requires changing the definition of the estimators $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$, while the second relaxes the assumption by studying the asymptotic distribution of the original estimators without the condition of a unique maximizer.
  • (A)
    Changing the estimators: The first alternative is to change the definition of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$ to avoid the optimization problem. For example, $t$ could be fixed at $t=0$ or learned from other folds inside the cross-fitting procedure. Focusing on the case of $\widehat{\theta}_{L}$, when estimating $\hat{s}_{L,k}$, one could estimate $\hat{t}_{k}$ (with data from all except the $k$-th fold) which maximizes the expression inside the $k$-th fold, and estimate the lower bound as
    $\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})$
    This would be valid since it is equivalent to a new estimator $\tilde{s}_{L,k}=\hat{s}_{L,k}+\hat{t}_{k}$, and there is no optimization in the final estimator.
  • (B)
    Asymptotic distribution without unique maximizer: For simplicity, I focus on the case of the lower bound. Define
    $\widehat{\Delta}(t)=\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t),$
    and note that $\widehat{\theta}_{L}=\max_{t}\widehat{\Delta}(t)$. From the proof of Theorem A.1 (eq. A.5), it follows that
    $\mathbb{G}_{n,k}(t)$ $=\sqrt{n}\left(\widehat{\Delta}(t,\hat{s}_{L,k})-\Delta_{P}(t,\hat{s}_{L,k})\right)$
    $=\sqrt{n}\left(\widehat{\Delta}(t,s_{L,P})-\Delta_{P}(t,s_{L,P})\right)+o_{P}(1)$
    converges to a Gaussian process. Note that $\sup_{t}\Delta_{P}(t,s_{L,P})\leq\theta^{*}_{L,P}$ (equation 2.3). If $\Delta_{P}(t,s_{L,P})\neq 0$ for at least one value of $t$, the $\sup$ function is Hadamard differentiable at $\Delta_{P}(\cdot,s_{L,P})$ (Cárcamo et al., 2020). It follows that
[…]
    uniformly in $P\in\mathcal{P}$, where $M=\{t:\Delta_{P}(t,s_{L,P})=\sup_{t^{\prime}}\Delta_{P}(t^{\prime},s_{L,P})\}$ and $\mathbb{G}_{P}$ is a Gaussian process with covariance function given by
    $\frac{Cov[Z_{t}(1),Z_{t^{\prime}}(1)]}{\pi}+\frac{Cov[Z_{t}(0),Z_{t^{\prime}}(0)]}{1-\pi},$
    where $Z_{t}(j)=\mathbb{I}(Y(j)\leq s_{L,P}+t)$, $\pi=P(D=1)$.
    Inference can be conducted by approximating the distribution of $\sup_{t\in M}\mathbb{G}_{P}(t)$. This could be done, for example, by conservatively estimating $M$ with $\widehat{M}$, such that $M\subseteq\widehat{M}$ with probability approaching one. Alternatively, one could apply the bootstrap. The validity of the bootstrap follows by combining the argument of Cárcamo et al. (2020), who show that the $\sup$ function is Hadamard differentiable, with a conditional functional delta method argument (e.g., Theorem 3.10.11 in Van Der Vaart and Wellner (2023)).
OA.4 Additional Tables